OpenStudy (anonymous):

Find the value of y in the sequence 2, 8, 3y +5,... 1. If the sequence was arithmetic 2. If the sequence was geometric

OpenStudy (anonymous):

1. common difference = 6 so next term = 14 3y+ 5 = 14 y = 3 2. common ratio = 8/2 = 4 so 3y + 5 = 32 3y = 27 y = 9